OpenStudy (anonymous):

multiplicative inverse of \[1\div3+i\]

OpenStudy (anonymous):

i could have sworn we just did this

OpenStudy (anonymous):

nope!

OpenStudy (anonymous):

\[\frac{1}{3+i}\]?

OpenStudy (anonymous):

multiplicative inverse is \[3+i\] because if you multiply them together you get 1

OpenStudy (anonymous):

nope!

OpenStudy (anonymous):

nope!

OpenStudy (anonymous):

so the answer is 3+i?

OpenStudy (anonymous):

it is true that \[\frac{1}{3+i}\times (3+i)=1\]

OpenStudy (anonymous):

complex numbers are weird, but not that weird

OpenStudy (anonymous):

hello joe hello fred

OpenStudy (anonymous):

i believe you satellite!

OpenStudy (owlfred):

xavier11, it's best when you put work in your questions so you understand how to derive the answer.

OpenStudy (anonymous):

nope!

OpenStudy (anonymous):

yes i completely agree, im just a little pressed for time right now